# EventFlow: Forecasting Continuous-Time Event Data with Flow Matching

### v1

@@ 4.3 Interpolant Construction
Informally, we view $v_{s}^{z}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ as prescribing a constant (but different) velocity to each of the $n$ events. For a fixed pair $z=(\gamma_{0},\gamma_{1})$ and a given sample $\gamma_{0}^{\prime}\sim\eta_{0}^{z}$, solving the system of ordinary differential equations $\,\mathrm{d}\gamma_{s}^{\prime}=v_{s}^{z}(\gamma_{s}^{\prime})\,\mathrm{d}s$ with initial condition $\gamma_{0}^{\prime}$ will result in a collection of events which is concentrated around the true event times $\gamma_{1}$. Note that here, we view this differential equation as an ODE in $\mathcal{T}^{n}$. If we draw many samples $\gamma_{0}\sim\eta_{0}^{z}$ and solve the corresponding ODE, the distribution over events at any intermediate time $s$ will be given by $\eta_{s}^{z}$.
In other words, the vector field $v_{s}^{z}$ generates the path of distributions $\eta_{s}^{z}$. However, this path is conditioned on $z$, and we would like to find the vector field $v_{s}$ which generates the unconditional path $\eta_{s}$. As is standard in flow matching (Lipman et al., 2023; Tong et al., 2024; Albergo & Vanden-Eijnden, 2023), the unconditional vector field $v_{s}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ may be obtained through a weighted marginalization procedure via
@@
4.4 Training, Parametrization, and Sampling
To train the model, we would like to perform regression on the vector fields $v_{s}$ in Equation (7). If we knew this vector field $v_{s}$, we could draw samples from the data TPP by drawing a sample event sequence $\gamma_{0}\sim\mu_{0}$ from the reference TPP, and flowing each event along the vector field $v_{s}$.
1 Sample $\gamma_{1}\sim\mu_{1}$, $s\sim\mathcal{U}[0,1]$, and $\epsilon\sim\mathcal{N}(0,1)$
@@ 2 $e_{\mathcal{H}}=\varnothing$
5       Sample split time $T_{0}\in[\Delta T,T-\Delta T]$
6       Construct history $\mathcal{H}\leftarrow\{t\in\gamma_{1}:t\leq T_{0}\}$
7       Embed history $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H})$
8       Construct future $\gamma_{1}\leftarrow\{t\in\gamma_{1}:T_{0}<t\leq T_{0}+\Delta T\}$
@@
10Set $n\leftarrow N(\gamma_{1})$
11 Sample $t_{0}^{1},\dots,t_{0}^{n}\sim q$ and sort to construct $\gamma_{0}$
12 Compute $\gamma_{s}^{z}$ via $t_{s}^{k}=(1-s)t_{0}^{k}+st_{s}^{k}$
@@
Algorithm 1 Training Step for EventFlow
Training
Foremost, although the marginal vector field in Equation (7) admits an analytical form, it is intractable to compute in practice as the marginal measure $\eta_{s}$ is not available. To overcome this, we may instead perform regression on the conditional vector fields $v_{s}^{z}$. Here, $v_{\theta}(\gamma_{s},s)$ will represent a neural network with parameters $\theta$ which takes in a sequence $\gamma_{s}$ of $N(\gamma_{s})=n$ event times, along with the flow time $s$. That is, we seek to minimize the loss
which can be shown to be equal to MSE regression on the unconditional $v_{s}$ up to an additive constant not depending on $\theta$ (Lipman et al., 2023; Tong et al., 2024). We note here that, although the regression target $v_{s}^{z}$ is linear, the unconditional vector field $v_{s}$ will in general be nonlinear. In practice, this loss is estimated by uniformly sampling a flow time $s\in[0,1]$, a pair $z=(\gamma_{0},\gamma_{1})\sim\rho$ from our balanced independent coupling, and drawing a noisy interpolant $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ according to Equation (5).
To train the model on a forecasting task, where the goal is to predict a future sequence of events conditioned on a history $\mathcal{H}$, we embed $\mathcal{H}$ into a fixed-dimensional vector representation $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$ via a learned encoder $f_{\theta}$ before providing this to the model $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$ and minimizing Equation (8). Note that we jointly train the encoder $f_{\theta}$ and vector field $v_{\theta}$. See Algorithm 1.
Parametrization
The second challenge is that we must learn a vector field $v_{\theta}(\gamma,s)$ in $n$ dimensions for arbitrary values of $n$. In other words, $v_{\theta}$ is a neural network which takes in a flow time $s\in[0,1]$ and a sequence of events $\gamma$, and must produce $N(\gamma)$ scalar outputs. We achieve this through an attention-based architecture, which we detail in Appendix C. At a high level, the flow time $s$ is transformed via a learnable embedding into a fixed-dimensional vector. Similarly, each event in $\gamma$ is transformed into fixed-dimensional vector via a learned embedding (which is shared across the events, but not the flow time). The flow-time embedding is then added to each event embedding, and the resulting sequence is passed through a standard transformer architecture (Yang et al., 2022; Vaswani, 2017), resulting in a sequence of $N(\gamma)$ vectors. Finally, each of these vectors is projected into one dimension via a linear layer to produce the sequence of $N(\gamma)$ velocities.
For conditional tasks, we must also compute an encoding $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$ of the history $\mathcal{H}$. This is done by a separate transformer encoder, which operates in the same fashion as described in the previous paragraph, but without the use of the flow-time $s$ as an input and without the final linear projection layer. This embedding is fed into the intermediate layers of our velocity network via cross-attention.
Lastly, for forecasting tasks we must also learn a model of the event count distribution $p_{\phi}(n\mid\mathcal{H})$. We treat this as a classification problem, where the goal is to predict the number of events $n$ occurring in the forecast window given the history $\mathcal{H}$. We again use an attention-based model, analogous to our velocity model, but where we aggregate the final sequence embedding by averaging and passing this through a small MLP. The model $p_{\phi}(n\mid\mathcal{H})$ is trained by minimizing the cross-entropy loss.
@@ 2 $e_{\mathcal{H}}=\varnothing$
5       Embed history $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H}$)
6       Sample $n\sim p_{\phi}(n\mid\mathcal{H})$
@@ 9       Sample $n\sim\mu_{1}(n)$
11Sample $t_{0}^{1},\dots,t_{0}^{n}\sim q$ and sort to construct $\gamma_{0}$
@@ 14       $\gamma_{s_{k}}\leftarrow\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1},e_{\mathcal{H}})$
15
Algorithm 2 Sampling Step for EventFlow
@@
Once the network $v_{\theta}$ is learned, we may draw samples from the model by first drawing a reference sequence $\gamma_{0}\sim\mu_{0}$ and solving the corresponding system of ODEs parametrized by $v_{\theta}$. More concretely, we first fix a number of events $n$ to generate. When seeking to unconditionally generate new sequences from the underlying data TPP $\mu_{1}$, we simply sample $n$ from the empirical event count distribution $\mu_{1}(n)$. For conditional tasks, we draw $n\sim q(n\mid\mathcal{H})$ from the learned conditional distribution over event counts.
Next, we draw $n$ initial events, corresponding to $s=0$, by sampling and sorting $t_{0}^{1},\dots,t_{0}^{n}\sim q$. In practice, we use $q=\mathcal{N}(0,I_{n})$ as we normalize our sequences into the range $[-1,1]$ during training and sampling (followed by renormalization to the data scale). Since we have fixed $n$, we may view this initial draw as a vector $\gamma_{0}=[t_{0}^{1},\dots,t_{0}^{n}]\in\mathcal{T}^{n}$. This event sequence $\gamma_{0}$ then serves as the initial condition for the system of ODEs $d\gamma_{s}=v_{\theta}(\gamma_{s},s)\,\mathrm{d}s$ which can be solved numerically. In our experiments, we use the forward Euler scheme, i.e., we specify a discretization $\{0=s_{0}<s_{1}<\dots<s_{K}=1\}$ of the flow time (in practice, uniform) and recursively compute
@@
where $h_{k}=s_{k}-s_{k-1}$ represents a scalar step size. While other choices of numerical solvers are certainly possible, we found that this simple scheme was sufficient as the model sample paths are typically close to linear. See Algorithm 2 for the full procedure.
@@ 5 Experiments
We study our proposed EventFlow model under two settings. The first is a conditional generative modeling task, where we seek to forecast both the number and times of future events given a history. The second is an unconditional task, where we aim to learn a representation of the underlying TPP distribution from empirical observations and generate new sequences from this distribution. In a sense, this second task can be viewed as a special case of the first with no observed history. Our overall experimental procedure is inspired by that of Lüdke et al. (2023).
Datasets
We evaluate our model across a diverse set of datasets encompassing a wide range of possible point process behaviors. First, we use a collection of six synthetic datasets produced by Omi et al. (2019). We additionally evaluate our model on seven real-world datasets, which are a standard benchmark for modeling unmarked TPPs (Shchur et al., 2020b; Bosser & Taieb, 2023; Lüdke et al., 2023). See Appendix A for additional information regarding our datasets.
@@
As a point of comparison, we train and evaluate several baseline models. These models were selected as they constitute a set of diverse and highly performant models. For an intensity-based method, we compare against the Neural Hawkes Process (NHP) (Mei & Eisner, 2017). We additionally compare against two intensity-free methods, namely the flow-based IFTPP model (Shchur et al., 2020a) and the diffusion-based model of Lin et al. (2022). Lastly, our strongest baseline is the recently proposed Add-and-Thin model of Lüdke et al. (2023), which can be loosely viewed as a non-autoregressive diffusion model. These models use an RNN-based history encoder, with the exception of Add-and-Thin which uses a CNN-based encoder. See Appendix D for additional details.
@@
Evaluating generative TPP models is challenging, as one must take into account both the variable locations and numbers of events. This is particularly challenging for the unconditional setting, where unlike forecasting, we do not have a ground-truth sequence to compare against. Our starting point is a metric (Xiao et al., 2017a) on the space of sequences $\Gamma$, allowing us to measure the distance between two sequences $\gamma=\sum_{k=1}^{n}\delta[t_{k}^{\gamma}]$ and $\eta=\sum_{k=1}^{m}\delta[t_{k}^{\eta}]$ with possibly different numbers of events. Without loss of generality, we assume $n\leq m$, so the distance is given by
$d(\gamma,\eta)=\sum_{k=1}^{n}|t_{k}^{\gamma}-t_{k}^{\eta}|+\sum_{k=n+1}^{m}(T-t_{k}^{\eta})$ (10)
where we recall that we assume the sequences are supported on $\mathcal{T}=[0,T]$. This distance can be understood either as an $L^{1}$ distance between the corresponding counting processes of $\gamma,\eta$ or as a generalization of the $1$-Wasserstein distance to measures of unequal mass. This metric allows us to compare two sequences of possibly unequal lengths, and will be used in our forecasting experiment.
For our unconditional experiment, we require a metric which will capture the distance between the TPP distributions themselves. To do so, we follow the approach suggested by Shchur et al. (2020b) and Lüdke et al. (2023), and use the distance in Equation (10) to calculate an MMD (Gretton et al., 2012). Namely, the MMD between two point process distributions $\mu,\nu\in\mathbb{P}(\Gamma)$ is given by
$\mathsf{MMD}(\mu,\nu)=\mathbb{E}_{\gamma,\gamma^{\prime}\sim\mu}[k(\gamma,\gamma^{\prime})]-2\mathbb{E}_{\gamma\sim\mu,\eta\sim\nu}[k(\gamma,\eta)]+\mathbb{E}_{\eta,\eta^{\prime}\sim\nu}[k(\eta,\eta^{\prime})]$ (11)
where $k:\Gamma\to\Gamma\to\mathbb{R}_{\geq 0}$ is a specified kernel. In accordance with prior work (Shchur et al., 2020b; Lüdke et al., 2023), we use an exponential kernel $k(\gamma,\eta)=\exp\left(-d(\gamma,\eta)/(2\sigma^{2})\right)$ with $\sigma$ chosen to be the median distance between all sequences.
@@ 5.1 Forecasting Event Sequences
Table 1: Sequence distance (10) between the forecasted and ground-truth event sequences on a held-out test set. Lower is better. We report the mean $\pm$ one standard deviation over five random seeds. The best mean distance on each dataset is indicated in bold, and the second best by an underline.
PUBG Reddit-C Reddit-S Taxi Twitter Yelp-A Yelp-M
IFTPP $2.6{\scriptstyle\pm 0.38}$ $20.4{\scriptstyle\pm 1.74}$ $43.8{\scriptstyle\pm 5.70}$ $3.4{\scriptstyle\pm 0.17}$ $1.9{\scriptstyle\pm 0.15}$ $1.4{\scriptstyle\pm 0.12}$ $2.6{\scriptstyle\pm 0.21}$
NHP $2.2{\scriptstyle\pm 0.07}$ $29.9{\scriptstyle\pm 2.17}$ $67.6{\scriptstyle\pm 1.2}$ $3.4{\scriptstyle\pm 0.50}$ $2.6{\scriptstyle\pm 0.53}$ $1.1{\scriptstyle\pm 0.04}$ $2.1{\scriptstyle\pm 0.52}$
Diffusion $4.5{\scriptstyle\pm 1.26}$ $20.8{\scriptstyle\pm 2.4}$ $64.5{\scriptstyle\pm 19.8}$ $3.8{\scriptstyle\pm 0.51}$ $1.8{\scriptstyle\pm 0.18}$ $1.1{\scriptstyle\pm 0.02}$ $2.6{\scriptstyle\pm 0.12}$
Add-and-Thin $\underline{2.0}{\scriptstyle\pm 0.03}$ $\underline{13.7}{\scriptstyle\pm 0.46}$ $\underline{21.5}{\scriptstyle\pm 0.51}$ $\underline{2.2}{\scriptstyle\pm 0.06}$ $\underline{1.5}{\scriptstyle\pm 0.06}$ $\underline{0.9}{\scriptstyle\pm 0.02}$ $\underline{1.4}{\scriptstyle\pm 0.04}$
EventFlow (ours) $\mathbf{0.5}{\scriptstyle\pm 0.01}$ $\mathbf{4.8}{\scriptstyle\pm 0.25}$ $\mathbf{6.8}{\scriptstyle\pm 0.44}$ $\mathbf{1.0}{\scriptstyle\pm 0.10}$ $\mathbf{0.4}{\scriptstyle\pm 0.02}$ $\mathbf{0.3}{\scriptstyle\pm 0.02}$ $\mathbf{0.5}{\scriptstyle\pm 0.04}$
We first evaluate our model on a multi-step conditional forecasting task, focusing on the real-world datasets. To do so, we set a forecast horizon $\Delta T$ for each of our real-world datasets, and generate event sequences in the range $[T_{0},T_{0}+\Delta T]$ for some given $T_{0}$, conditioned on the history of events $\mathcal{H}_{T_{0}}$. Up to a shift, this means we are taking $\mathcal{T}=[0,\Delta T]$ as the support for our model TPP. The forecast horizon $\Delta T$ is chosen such that the window typically contains multiple events.
At training time, we uniformly sample $T_{0}\in[\Delta T,T-\Delta T]$ and split a given data sequence $\gamma_{1}$ into a history on $[0,T_{0}]$ and a future $[T_{0},T_{0}+\Delta T]$. We encode the history $\mathcal{H}_{T_{0}}$ before training the model to fit the events occurring in the future. At testing time, we perform the same splitting procedure, sampling 50 values of $T_{0}$ for each test set sequence. We then forecast the sequence in $[T_{0},T_{0}+\Delta T]$ via the model and compute the distance (10) between the ground-truth and generated sequences.
@@
We report the results of this experiment in Table 1. Our proposed EventFlow method obtains the lowest average forecasting error across all datasets, obtaining an error which is 2.2 to 4 times lower than the second-best method, Add-and-Thin. Given that the non-autoregressive models (EventFlow, Add-and-Thin) consistently outperform the autoregressive baselines, this is clear evidence that autoregressive models can struggle on multi-step predictions. This is especially true on the Reddit-C and Reddit-S datasets, which exhibit long sequence lengths.

### v2

@@ 4 EventFlow
In other words, flowing samples $\hat{\gamma}_{0}^{z}$ along the vector field $v_{s}^{z}$ generates the path of distributions $\eta_{s}^{z}$. However, this path is conditioned on $z$, and we would like to find the vector field $v_{s}$ which generates the unconditional path $\eta_{s}$. As is standard in flow matching (Lipman et al., 2023; Tong et al., 2024; Albergo & Vanden-Eijnden, 2023), the unconditional vector field $v_{s}:\mathcal{T}^{n}\to\mathbb{R}^{n}$ may be obtained via
@@
1 Sample $\gamma_{1}\sim\mu_{1}$, $s\sim\mathcal{U}[0,1]$, $\epsilon\sim\mathcal{N}(0,1)$
@@ 2 $e_{\mathcal{H}}=\varnothing$
5       Sample $T_{0}\in[\Delta T,T-\Delta T]$
6       $\mathcal{H}\leftarrow\{t\in\gamma_{1}:t\leq T_{0}\}$
7       $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H})$
8       $\gamma_{1}\leftarrow\{t\in\gamma_{1}:T_{0}<t\leq T_{0}+\Delta T\}$
@@
[…]
11 Sample and sort $t_{0}^{k}\sim q$ for $k=1,\dots,n$
12 $\gamma_{0}\leftarrow(t_{0}^{1},\dots,t_{0}^{n})$
13 $t_{s}^{k}\leftarrow(1-s)t_{0}^{k}+st_{1}^{k}$ for $k=1,\dots,n$
14 $\gamma_{s}^{z}\leftarrow(t_{s}^{1},\dots,t_{s}^{n})$
@@
Algorithm 1 Training EventFlow
@@ 2 $e_{\mathcal{H}}=\varnothing$
5       $e_{\mathcal{H}}\leftarrow f_{\theta}(\mathcal{H}$)
6       $n\sim p_{\phi}(n\mid\mathcal{H})$
@@ 9       $n\sim\mu_{1}(n)$
11Sample and sort $t_{0}^{k}\sim q$ for $k=1,\dots,n$
12 $\gamma_{0}\leftarrow(t_{0}^{1},\dots,t_{0}^{n})$
@@ 15       $\gamma_{s_{k}}\leftarrow\gamma_{s_{k-1}}+h_{k}v_{\theta}(\gamma_{s_{k-1}},s_{k-1},e_{\mathcal{H}})$
16
return $\gamma_{0}$
Algorithm 2 Sampling EventFlow
Training
If we knew $v_{s}$ in Equation (6), we could draw samples from $\mu_{1}$ by drawing a sample event sequence $\gamma_{0}\sim\mu_{0}$ from the reference TPP and flowing each event along $v_{s}$. Although the marginal vector field in Equation (6) admits an analytical form, it is intractable to compute in practice as the marginal measure $\eta_{s}$ is not available. To overcome this, we instead regress on the conditional vector fields $v_{s}^{z}$. Here, $v_{\theta}(\gamma_{s},s)$ will represent a neural network with parameters $\theta$ which takes in a sequence $\gamma_{s}$ along with the flow time $s$. We seek to minimize the loss
which previous works on flow matching have shown to be equal to MSE regression on the unconditional $v_{s}$ up to an additive constant not depending on $\theta$ (Lipman et al., 2023; Tong et al., 2024). We estimate $J(\theta)$ by uniformly sampling a flow time $s\in[0,1]$, a pair $z=(\gamma_{0},\gamma_{1})\sim\rho$ from our balanced coupling and drawing a noisy interpolant $\hat{\gamma}_{s}^{z}\sim\eta_{s}^{z}$ according to Equation (4).
To train the model on a forecasting task, where the goal is to predict a future sequence of events conditioned on a history $\mathcal{H}$, we embed $\mathcal{H}$ into a fixed-dimensional vector representation $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$ via a learned encoder $f_{\theta}$ before providing this to the model $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$ and minimizing Equation (7). Note that we jointly train the encoder $f_{\theta}$ and vector field $v_{\theta}$. In this case, we must also learn a model of the event count distribution $p_{\phi}(n\mid\mathcal{H})$. We train $p_{\phi}(n\mid\mathcal{H})$ by minimizing a cross-entropy loss. In practice, we found it important to regularize this loss to encourage smoothness in $n$, which we achieve through an optimal-transport based regularizer. As both $v_{\theta}$ and $p_{\phi}$ must be able to operate on variable-length sequences, we use a transformer-based backbone for both models. However, our overall method is agnostic to this choice. See Algorithm 1 and Appendix C.
@@
Once $v_{\theta}$ is learned, we may sample from the model by drawing a reference sequence $\gamma_{0}\sim\mu_{0}$ and solving the corresponding ODE parametrized by $v_{\theta}$. More concretely, we first fix a number of events $n$. When seeking to unconditionally generate new sequences from the underlying data TPP $\mu_{1}$, we simply sample $n$ from the empirical event count distribution $\mu_{1}(n)$. For conditional tasks, we draw $n\sim p_{\phi}(n\mid\mathcal{H})$ from the learned conditional distribution over event counts. Next, we draw $n$ initial events, corresponding to $s=0$, by sampling and sorting $t_{0}^{1},\dots,t_{0}^{n}\sim q$. In practice, we use $q=\mathcal{N}(0,I_{n})$ as we normalize our sequences into the range $[-1,1]$ during training and sampling (followed by renormalization to the data scale). Since we have fixed $n$, we may view this initial draw as a vector $\gamma_{0}=[t_{0}^{1},\dots,t_{0}^{n}]\in\mathcal{T}^{n}$. This event sequence $\gamma_{0}$ then serves as the initial condition for the system of ODEs $d\gamma_{s}=v_{\theta}(\gamma_{s},s)\,\mathrm{d}s$ which can be solved numerically. In our experiments, we use the forward Euler scheme, i.e., we specify a discretization $\{0=s_{0}<s_{1}<\dots<s_{K}=1\}$ of the flow time (in practice, uniform) and recursively compute
@@
where $h_{k}=s_{k}-s_{k-1}$ is a step size. While other choices are possible, we found that this simple scheme was sufficient as the model sample paths are typically close to linear. See Algorithm 2.
@@ 5 Experiments
We study our proposed EventFlow model under two settings. The first is a conditional task, where we seek to forecast both the number of future events and their times of occurrence, over a particular horizon, given a history. The second is an unconditional task, where we aim to learn a representation of the underlying TPP distribution from empirical observations and generate new sequences from this distribution. This second task can be viewed as a special case of the first with no observed history. We evaluate our model across a diverse set of datasets encompassing a wide range of possible point process behaviors. We evaluate our model on seven real-world datasets, which are a standard benchmark for modeling unmarked TPPs (Shchur et al., 2020b; Bosser & Taieb, 2023; Lüdke et al., 2023). We also evaluate on a collection of six synthetic datasets (Omi et al., 2019). See Appendix A.
@@
We selected a set of baselines consisting of a set of diverse and highly performant neural TPP models. For an intensity-based method, we compare against the Neural Hawkes Process (NHP) (Mei & Eisner, 2017), a well-known and very widely-used neural TPP model. We additionally compare against two intensity-free methods, namely the flow-based IFTPP model (Shchur et al., 2020a) and the diffusion-based model of Lin et al. (2022). Lastly, our strongest baseline is the recently proposed Add-and-Thin model of Lüdke et al. (2023), which can be loosely viewed as a non-autoregressive diffusion model. These models use an RNN-based history encoder, with the exception of Add-and-Thin which uses a CNN-based encoder. See Appendix D for additional details.
@@
Evaluating generative TPP models is challenging, as one must take into account both the variable locations and numbers of events. This is particularly challenging for the unconditional setting, where unlike forecasting, we do not have a ground-truth sequence to compare against. Our starting point is a metric (Xiao et al., 2017a) on the space of sequences $\Gamma$, allowing us to measure the distance between two sequences $\gamma=\sum_{k=1}^{n}\delta[t^{k}_{\gamma}]$ and $\eta=\sum_{k=1}^{m}\delta[t^{k}_{\eta}]$ with possibly different numbers of events. Without loss of generality, we assume $n\leq m$, where the distance is given by
$d(\gamma,\eta)=\sum_{k=1}^{n}|t^{k}_{\gamma}-t^{k}_{\eta}|+\sum_{k=n+1}^{m}(T-t^{k}_{\eta})$ (9)
where we recall that sequences are supported on $\mathcal{T}=[0,T]$. This distance can be understood either as an $L^{1}$ distance between the counting processes of $\gamma,\eta$ or as a generalization of the $1$-Wasserstein distance to measures of unequal mass. For our unconditional experiment, we require a metric which will capture the distance between the TPP distributions themselves. To do so we use the distance in Equation (9) to calculate an MMD (Gretton et al., 2012; Shchur et al., 2020b). We use an exponential kernel $k(\gamma,\eta)=\exp\left(-d(\gamma,\eta)/(2\sigma^{2})\right)$ with $\sigma$ chosen to be the median distance between all sequences (Shchur et al., 2020b; Lüdke et al., 2023).
Figure 2: Sequence distance (9) between the forecasted and ground-truth event sequences on a held-out test set. We report the mean $\pm$ one standard deviation over five random seeds. EventFlow (with 25 NFEs) achieves the lowest mean distance (forecasting error) for each of the 7 datasets.
@@ 5.1 Forecasting Event Sequences
We first evaluate our model on a multi-step forecasting task. We set a horizon $\Delta T$ for each of our real-world datasets and seek to generate event sequences in the range $[T_{0},T_{0}+\Delta T]$ for some given $T_{0}$, conditioned on the history $\mathcal{H}_{T_{0}}$. Up to a shift, this means we are taking $\mathcal{T}=[0,\Delta T]$. At training time, we uniformly sample $T_{0}\in[\Delta T,T-\Delta T]$. At test time, we sample 50 values of $T_{0}$ for each test set sequence. We then forecast the sequence in $[T_{0},T_{0}+\Delta T]$ and compute the distance (9) between the ground-truth and generated sequences. We note that the distance in Equation (9) is computed using $T_{0}+\Delta T$ rather than $T$ as the maximum event time, as using $T$ would result in a distance which is sensitive to the location of the forecasting window. We further normalize Equation (9) by $\Delta T$. This overall approach is similar to that of Lüdke et al. (2023). For all datasets, the sequences are split into train-validation-test sets, and all metrics are reported on the held-out test set.
@@
We report the results of these experiments in Figure 2. Our proposed EventFlow method obtains the lowest average forecasting error across all datasets, achieving an error which is 20%-53% lower than the nearest baseline. Given that the non-autoregressive models (EventFlow, Add-and-Thin) consistently outperform the autoregressive baselines, this is clear evidence that autoregressive models can struggle on multi-step predictions. This is especially true on the Reddit-C and Reddit-S datasets, which exhibit long sequence lengths.
Table 1: Mean sequence distance per dataset for EventFlow as we vary the NFEs used to simulate the ODE.
NFEs PUBG Red.C Red.S Taxi Twitter Yelp-A Yelp-M
25 $2.0$ $15.8$ $16.0$ 3.23.2{3.2}3.2 1.41.4{1.4}1.4 1.31.3{1.3}1.3 1.91.9{1.9}1.9
10 $2.0$ $15.8$ $15.8$ $3.1$ $1.4$ $1.3$ $1.9$
1 $2.0$ $15.8$ $15.8$ $3.7$ $1.4$ $1.8$ $1.9$
In Table 1, we additionally vary the number of function evaluations (NFEs) used at test time, i.e., the number of steps $K$ used in the ODE solver in Algorithm 2. We find that reducing $K$ results in minimal (or no) loss in performance. Even with only a single NFE, EventFlow is able to obtain state-of-the-art performance. This is because, although the vector fields (6) in EventFlow are generally non-linear, we find that our construction produces paths which are approximately so. In contrast, Add-and-Thin uses 100 NFEs, while the NFEs for the autoregressive models scale linearly with the number of sampled events.
Table 2: MARE values evaluating the predicted number of events when forecasting. Mean values are reported over five random seeds.
PUBG Red.C Red.S Taxi Twitter Yelp-A Yelp-M
IFTPP $1.05$ $1.69$ $0.79$ $0.60$ $0.88$ $0.76$ $0.76$
NHP $1.02$ $\underline{0.95}$ $1.00$ $0.67$ $2.48$ $0.80$ $1.07$
Diffusion $1.95$ $1.28$ $1.12$ $0.49$ $0.66$ $0.65$ $0.72$
Add-and-Thin $\underline{0.43}$ $0.99$ $\underline{0.38}$ $\underline{0.33}$ $\underline{0.60}$ $\mathbf{0.42}$ $\mathbf{0.46}$
EventFlow (ours) $\mathbf{0.40}$ $\mathbf{0.70}$ $\mathbf{0.16}$ $\mathbf{0.28}$ $\mathbf{0.46}$ $\underline{0.56}$ $\underline{0.50}$
To evaluate the predictions of event counts, ${p_{\phi}(n\mid\mathcal{H})}$, we report the mean absolute relative error (MARE) between the true and predicted event counts. This is reported in Table 2. Our model achieves strong performance on this metric as it decouples the event count prediction from the event time prediction, allowing us to train explicitly for each task. The baselines, on the other hand, only model $n$ implicitly. See Appendix E for details and standard deviations.